A piece of poster board has a perimeter of 156 in. If Mika cut 6 in off the width, the length becomes four times the width. What are the dimensions of the poster board?

Answer by TimothyLamb(4379)

p = 2L + 2w = 156
---
L = 4(w - 6)
---
2L + 2w = 156
2(4(w - 6)) + 2w = 156
8w - 48 + 2w = 156
10w = 204
---
w = 20.4 in
L = 57.6 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php